Solve -3*e^(2x)+7=120
Answer by Alan3354(69443)

Solve -3*e^(2x)+7=120
-3e^(2x) = 113
e^(2x) = -113/3
2x = ln(-113/3)
logs of negative numbers not allowed.
--> no solution